OpenStudy (anonymous):

A rectangle has a length of 5.50 m and a width of 12.0 m. What are the perimeter and area of this rectangle?

OpenStudy (anonymous):

A= l * w A = 5.50m * 12.0m= 66 sq m P= l+l+w=w P 5.50+5.50+12.0+12.0 =35m